OpenStudy (anonymous):

If the perimeter of a rectangle is 68 inches and the length is 12 inches longer than the width, what is the width of the rectangle?

OpenStudy (ivettef365):

P = 2l + 2w l = w+12 P = 68

OpenStudy (ivettef365):

now just substitute as follows: 68 = 2(w+12) + 2w (solve for w)

OpenStudy (anonymous):

wat is it

OpenStudy (ivettef365):

do you know how to solve the equation ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its 11 : )

OpenStudy (anonymous):

thanks

OpenStudy (ivettef365):

good job